Below we present data concerning hospital labor needs at 13 U.S. Navy Hospitals. Here: y monthly labor hours required and x =Consider the data in the information page: The value of SSE is: (Leave the answer as a whole number) Answer
Answer #1
Σn ΣX ΣY ΣXY ΣX² ΣY²
13 26517.68 28388.79 74559098.00 72900841.00 79037513.00

              
sample size ,   n =   13      
here, x̅ =Σx/n =   2039.8215   ,   ȳ = Σy/n =   2183.753077
              
SSxx =    Σx² - (Σx)²/n =   18809506.186      
SSxy=   Σxy - (Σx*Σy)/n =   16651032.707      
SSyy =    Σy²-(Σy)²/n =   17043405.487      

SSE=   (Sx*Sy - S²xy)/Sx = (18809506.186*17043405.487 - 16651032.707²)/18809506.186

= 2303151.9436 ~ 2303152 (answer)